OpenStudy (anonymous):

HELP please? :) a line includes points 0 comma negative 1 and 5 comma negative 2. A line includes points negative 5 comma negative 8 and 0 comma negative 5 3x + 5y = –25 x – 5y = 5 3x + 5y = –25 x + 5y = –5 3x – 5y = 25 x + 5y = –5 3x – 5y = 25 x – 5y = 5

OpenStudy (anonymous):

Those are the choses...hold on ill try and get the picture.

OpenStudy (anonymous):

I'm not sure how to post it here..

OpenStudy (waleed_imtiaz):

3x + 5y = –25 x + 5y = –5 these are the correct equations @magenxoxo ... use two point form to solve the points given...

OpenStudy (anonymous):

You really wouldn't have to show any picture for this. Your problem description is sufficient. You have 2 pairs of points, where each pair determines a line, so you have 2 lines. For each pair, separately, you get the slope, m, by the slope formula: \[m = \frac{ y _{1} - y _{2} }{ x _{1} - x _{2} }\]where your points are of the form:\[(x _{1}, y _{1}) and (x _{2}, y _{2})\]Then you use the point-slope form for the equation of a line:\[y - y _{1} = m(x - x _{1})\]and you really at that point could use either point. Then you convert this point-slope equation of a line into standard form. And a note to the helpers, please let the students try to solve the problem, don't just give answers.

OpenStudy (anonymous):

Thank you waleed for the answer! but thank you Tcarr for explaining!!! :)))

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition, @magenxoxo

OpenStudy (waleed_imtiaz):

pleasure is always mine... !!! but try to do it urself using two point form ...points are given... and just check the answer... !

OpenStudy (anonymous):

@magenxoxo , do you want to go through it in a little more detail? Can you find the slopes ok?

OpenStudy (anonymous):

Sure, why not :)

OpenStudy (anonymous):

tcarroll010 will tell you many things.

OpenStudy (anonymous):

ok, at the higher conceptual level, you have 2 lines, each of which is determined by a pair of points, so you will be coming up with a slope for the first equation and another slope for the second equation. To start, you might want to plug in the points in that slope formula I gave you. You can either type the work or the final slope answers. and I'll check that for you.

OpenStudy (anonymous):

:-) hopefully some of those things will be "good things" :-) @chamel0305

OpenStudy (anonymous):

yeah. lol

OpenStudy (anonymous):

Here's a hint to get started with the slope of the first line. You can use use: x1 = 0, y1 = -1, x2 = 5, y2 = -2 because that lines up the form of: (x1, y1), (x2, y2) That's for the first pair of points and that is all you need to get the first slope. Once you get that, we can attempt the point-slope form of the equation of a line.

OpenStudy (anonymous):

Ok. 1 sec. :)

OpenStudy (anonymous):

Actually, I don't have time right now to work on this. I need to finish my other schol because im leaving soon. But ill keep this question open and come back to it :) THANK YOU FOR YOU'RE HELP!!!

OpenStudy (anonymous):

anytime! It was nice talking with you! You're very welcome!